OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2. (1/2A+1/2B)^2 a. Simplify b. What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

I'm not sure what part (a) is about. But if A represents P(event A) and similarly for B, you have (1/2 * 1/2 + 1/2 * 1/2)^2 = (1/4 + 1/4)^2 = (1/2)^2 = 1/4 If that's not what A and B represent, you have, more generally, (A/2 + B/2)^2 = A^2/4 + B^2/4 + 2 AB / 4 = A^2 / 4 + B^2 / 4 + AB / 2 For part (b): Roll one of the cubes. The top facing number will be either from {1 5 6} or {2 3 4} - it doesn't matter which. Then you want the other one to have its top face come from the other trio of numbers, so P(success) = 3/6 = 1/2 You could work out the formula with a longer computation: P(156 & 234) + P(234 & 156) = 2 * (1/2 * 1/2) = 2 * 1/4 = 1/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np